You have a rectangular space in your yard that you want to put a border around. You find that 60 feet of border material is needed. The length of the rectangle is exactly one-half the width. What are the dimensions of the space in your yard?
Answer by jessica43(140)

To solve this problem you need to use the given information to write two equations and then solve.
First, you know that a rectangle has 4 sides with each of the opposite sides having the same length, and the total perimeter is 60 feet. So:
W+W+L+L = 60 (with W=width and L= length)
or 2W + 2L = 60
Secondly, we know that the length is exactly one-half the width, so:
L = 1/2W
Now you can plug the second equation into the first equation, replacing the L value:
2W + 2L = 60
2W + 2(1/2W) = 60
2W + W = 60
3W = 60
W = 20
Now plug this into the second equation to find the length:
L = 1/2W
L = 1/2(20)
L = 10
So the length of the rectangle is 10 feet and the width is 20 feet.